OpenStudy (anonymous):

If a line with a slope of -2 crosses the y-axis at (0, 3), what is the equation of the line?

OpenStudy (anonymous):

y − 2x = 3 -2y + 3x = 1 y + 3x = 2 y + 2x = 3

OpenStudy (igreen):

D

OpenStudy (igreen):

y + 2x = 3 can be re-written as y = -2x + 3 It's in the form of: y = mx + b Where m = slope, and b = y-intercept. y = -2x + 3 Where -2 = slope and 3 = y-intercept.

OpenStudy (anonymous):

Find the slope and the y-intercept of the equation 2x − 3y = 6

OpenStudy (igreen):

y = -2/3x - 2 Slope = -2/3 and y-intercept = -2

OpenStudy (igreen):

@usitdown